# Loop-Residual Neural Networks for Iterative Refinement

### v1

@@ 3 Methodology
The Loop-Residual model introduces an iterative mechanism where a subset of layers or components is looped through multiple times for iterative refinement. Specifically, instead of passing the input through all layers sequentially, the model loops over a part of the network, predicting the residual between the current state and the desired state at each iteration.
@@ 3.1 General Loop Structure with Residual Design
Our core design involves refining the hidden state by iteratively looping over transformer blocks with residual connections. The process is defined as:
$x^{(n)}=x^{(n-1)}+f_{\theta}\left(x^{(n-1)}\right)$ (1)
@@
  • $x^{(n)}$ is the hidden state at iteration $n$.
  • $x^{(0)}$ is the initial hidden state from the previous layer or embedding layer.
  • $f_{\theta}$ represents the function parameterized by $\theta$, consisting of the looping transformer blocks predicting the residual.
At each iteration, the function $f_{\theta}$ predicts the residual error based on the current hidden state $x^{(n-1)}$, and this residual is added back to update the hidden state. This formulation allows the model to progressively refine its predictions by focusing on the residuals, enhancing its capacity to capture complex patterns without increasing the model size.
@@
Figure 1: Loop-Residual Mechanism: At each iteration, the transformer block(s) predict the residual, which is added back to the current hidden state.

### v2

@@ 3 Methodology
We propose a novel architecture called the Loop-Residual Neural Network, which refines the hidden state through iterative loops over a subset of transformer blocks with residual connections and gating mechanisms. This iterative process allows the model to revisit and enhance its internal representations multiple times, effectively increasing its approximation capacity without adding more parameters.
@@ 3.1 General Loop Structure with Residual Design
In our model, the hidden state is updated iteratively using an approach inspired by series expansions in numerical methods. Let $x^{(n)}\in\mathbb{R}^{d}$ denote the hidden state after the $n$-th iteration, where $x^{(0)}$ is the initial hidden state obtained from the embedding layer or the output of the previous layer.
The iterative update rule is defined as:
$x^{(n)}=x^{(n-1)}+a_{n}\odot f_{\theta}\left(x^{(n-1)}\right),$ (1)
@@
  • $x^{(n-1)}\in\mathbb{R}^{d}$ is the hidden state from the previous iteration.
  • $f_{\theta}\left(x^{(n-1)}\right)\in\mathbb{R}^{d}$ is the output of the transformer block(s) parameterized by $\theta$.
  • $a_{n}\in\mathbb{R}^{d}$ is a gating coefficient vector at iteration $n$, applied element-wise (denoted by $\odot$). The gating coefficients are learned parameters.
  • $a_{0}=\mathbf{1}$ (a vector of ones).
By unrolling the iterative process, the hidden state after $N$ iterations can be expressed as:
$x^{(N)}=a_{0}\odot x^{(0)}+\sum_{k=1}^{N}a_{k}\odot f_{\theta}\left(x^{(k-1)}\right).$ (2)
Alternatively, we can represent the hidden state accumulation explicitly:
$x^{(N)}=a_{0}\odot x^{(0)}+a_{1}\odot f_{\theta}\left(x^{(0)}\right)+a_{2}\odot f_{\theta}\left(x^{(1)}\right)+\dots+a_{N}\odot f_{\theta}\left(x^{(N-1)}\right).$ (3)
Since $a_{0}=\mathbf{1}$ and $f_{\theta}\left(x^{(k-1)}\right)$ represents the correction at each iteration, the model effectively accumulates the contributions from all iterations to refine the hidden state.
The iterative refinement process allows the model to capture complex patterns and dependencies by progressively updating the hidden state. The use of residual connections helps to alleviate issues such as vanishing gradients and facilitates the flow of information across iterations.
3.2 Illustration of the Loop-Residual Mechanism
@@
Figure 1: Computational Graph with Multiple Transformer Blocks